Two blocks of mass 3.50 kg and 8.00 kg are connected by a massless string that passes
over a frictionless pulley (Fig. below).The inclines are frictionless. Find (a) the
magnitude of the acceleration of each block and (b) the tension in the string
1
Expert's answer
2021-08-27T14:39:32-0400

a)


"-3.5(9.8)\sin 35+T=3.5a\\8(9.8)\sin 35-T=8a\\a=2.2\frac{m}{s^2}"

b)


"T=27.4\ N"